OpenStudy (gabylovesyou):

Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = 1/4; y intercept at (0, 1) slope = 1/4; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

jimthompson5910 (jim_thompson5910):

What do you get when you solve for y?

OpenStudy (gabylovesyou):

mmmm

jimthompson5910 (jim_thompson5910):

any ideas? or did you get anywhere with solving for y?

OpenStudy (gabylovesyou):

y = n - 1

jimthompson5910 (jim_thompson5910):

close, but you're dealing with x and when you divide everything by 4, you have to divide x by 4 too

OpenStudy (gabylovesyou):

y = x-4/4

jimthompson5910 (jim_thompson5910):

x - 4y = 4 x = 4+4y x-4 = 4y 4y = x - 4 y = x/4 - 4/4 y = 1/4 x - 1

jimthompson5910 (jim_thompson5910):

So what are the slope and y-intercept?

OpenStudy (gabylovesyou):

ohhh sooo its C?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (gabylovesyou):

Thank you soo much :)

jimthompson5910 (jim_thompson5910):

np